# Data-Driven Dynamic Factor Modeling via Manifold Learning

### v1

@@ 3.3 Joint Diffusion Kalman Filter
We describe here our linear state space approximation for the dynamics of the diffusion coordinates. Assuming that the gradients in (3.8) are constant, [shnitzer2020diffusion] define a state-space model, dubbed the Diffusion Kalman Filter (DKF), for the case where the data arises from a single data set $(x(t_{i}))_{i=0}^{N}$ (see Appendix A). We now modify DKF for a supervised learning framework to incorporate information in the series of the dependent variables $y(t_{i})$, which depend jointly with $x(t_{i})$ on $\theta(t_{i})$. If one were to use the estimated diffusion coordinates from DKF as developed in [shnitzer2020diffusion], a simple OLS regression of the dependent variables on these coordinates will show that most of the explanatory power of the original covariates $x(t_{i})$ for the dependent variables is lost (see, for instance, Appendix B for the regression of individual stock returns on a high-dimensional set of macroeconomic covariates). Thus, it is important to extend DKF to also account for the variation in the dependent variable in order for the diffusion coordinates to preserve explanatory power. We call this approach the Joint Diffusion Kalman Filter (JDKF). In this supervised learning framework, we adopt the dynamical system setting from Subsection 3.1:
@@
Following the procedure in Subsection 3.2, the diffusion coordinates $\psi(t_{i})$ are computed jointly on samples of $(X,Y)$, denoted $(x(t_{i}),y(t_{i}))_{i=0}^{N}$, and $x$ and $y$ can be similarly reconstructed by the linear lifts:
@@
where $\mathbf{H}_{i,j}^{x}=\langle x_{i},\psi_{j}\rangle_{\text{data}}$ and $\mathbf{H}_{i,j}^{y}=\langle y_{i},\psi_{j}\rangle_{\text{data}}$. Let $\Lambda$ denote the transition matrix for the linear dynamics containing the mapped eigenvalues ${-\varepsilon^{-1}\log\kappa_{1},\dots-\varepsilon^{-1}\log\kappa_{\ell}}$ on its diagonal.
@@
3.3.1 E-Step: Kalman Filter with Joint Observations
The Kalman filter is used to estimate the latent state $\psi_{t}$ and the corresponding state covariance $P_{t|t}$. In our supervised learning framework, we define the joint observation vector:
$z_{t}=\begin{bmatrix}x_{t}\\ y_{t}\end{bmatrix}=\begin{bmatrix}\mathbf{H}^{x}\\ \mathbf{H}^{y}\end{bmatrix}\psi_{t}+\begin{bmatrix}v_{x,t}\\ v_{y,t}\end{bmatrix}$
Note that in the linear case, we would replace $\mathbf{H}^{y}$ with $\mathbf{B}\mathbf{H}^{x}$. The steps of the Kalman filter are as follows:
  1. 1.
    Prediction Step:
    $\hat{\psi}_{t|t-1}=\mathbf{A}\hat{\psi}_{t-1|t-1},\quad P_{t|t-1}=\mathbf{A}P_{t-1|t-1}\mathbf{A}^{\top}+\mathbf{Q}$
  2. 2.
    Innovation (Residual):
    $e_{t}=z_{t}-\mathbf{H}^{z}\hat{\psi}_{t|t-1}$
    where:
    $\mathbf{H}^{z}=\begin{bmatrix}\mathbf{H}^{x}\\ \mathbf{H}^{y}\end{bmatrix}$
    and the joint noise covariance is:
    $\mathbf{R}=\begin{bmatrix}\mathbf{R}_{x}&\mathbf{R}_{xy}\\ \mathbf{R}^{\top}_{xy}&\mathbf{R}_{y}\end{bmatrix}$
  3. 3.
    Innovation Covariance:
    $S_{t}=\mathbf{H}^{z}P_{t|t-1}{\mathbf{H}^{z}}^{\top}+\mathbf{R}$
  4. 4.
    Kalman Gain:
    $K_{t}=P_{t|t-1}{\mathbf{H}^{z}}^{\top}S_{t}^{-1}$
  5. 5.
    Update Step:
    $\hat{\psi}_{t|t}=\hat{\psi}_{t|t-1}+K_{t}e_{t},\quad P_{t|t}=(I-K_{t}{\mathbf{H}^{z}}^{\top})P_{t|t-1}$
  6. 6.
    Log-Likelihood Contribution:
    $\mathcal{L}^{(t)}=-\frac{1}{2}\left(\log|S_{t}|+e_{t}^{\top}S_{t}^{-1}e_{t}+(n+m)\log 2\pi\right)$
After running the Kalman filter, we apply the Rauch–Tung–Striebel (RTS) smoother [rauch_tung_striebel]:
  1. 1.
    Initialize at $T$:
    $\hat{\psi}_{T|T}=\hat{\psi}_{T|T},\quad P_{T|T}=P_{T|T}$
  2. 2.
    Backward pass for $t=T-1,...,1$:
    $J_{t}=P_{t|t}\mathbf{A}^{\top}P_{t+1|t}^{-1}$
    $P_{t|T}=P_{t|t}+J_{t}(P_{t+1|T}-P_{t+1|t})J_{t}^{\top}$
3.3.2 M-Step: Parameter Updates
After running the Kalman filter, we update the parameters $\mathbf{R}_{x}$, $\mathbf{R}_{y}$, and $\mathbf{R}_{xy}$ (and $\mathbf{B}$ in the linear case) based on the current estimates of $\psi_{t}$.
  1. 1.
    Update $\mathbf{R}_{x}$ using the residuals from the smoothed estimates $\hat{\psi}_{t|T}$:
    $\mathbf{R}_{x}=\frac{1}{T}\sum_{t=1}^{T}\left(x_{t}-\mathbf{H}^{x}\hat{\psi}_{t}\right)\left(x_{t}-\mathbf{H}^{x}\hat{\psi}_{t}\right)^{\top}$
  2. 2.
    Update $\mathbf{R}_{y}$ using the residuals from the smoothed estimates $\hat{\psi}_{t|T}$:
    $\mathbf{R}_{y}=\frac{1}{T}\sum_{t=1}^{T}\left(y_{t}-\mathbf{H}^{y}\hat{\psi}_{t}\right)\left(y_{t}-\mathbf{H}^{y}\hat{\psi}_{t}\right)^{\top}$
  3. 3.
    Update $\mathbf{R}_{xy}$ using the residuals from the smoothed estimates $\hat{\psi}_{t|T}$:
    $\mathbf{R}_{xy}=\frac{1}{T}\sum_{t=1}^{T}\left(x_{t}-\mathbf{H}^{x}\hat{\psi}_{t}\right)\left(y_{t}-\mathbf{H}^{y}\hat{\psi}_{t}\right)^{\top}$
  4. 4.
@@ 3.3.2 M-Step: Parameter Updates
    In the linear case, update $\mathbf{B}$ via least squares regression on the lifted smoothed estimates $\hat{\psi}_{t|T}$:
    $\mathbf{B}=\left(\sum_{t=1}^{T}y_{t}(\mathbf{H}^{x}\hat{\psi}_{t})^{\top}\right)\left(\sum_{t=1}^{T}\mathbf{H}^{x}\hat{\psi}_{t}(\mathbf{H}^{x}\hat{\psi}_{t})^{\top}\right)^{-1}$
3.3.3 EM Algorithm
@@
We have added an extra layer with the measurement equation for the dependent variables to maintain the explanatory power in the embeddings and taken this into account via the modified state-space. JDKF thus gives us diffusion coordinates that propagate linearly and preserve much of the original explanatory power from the full high-dimensional set of covariates. Figure 3 displays a broad flow map of the JDKF procedure. In the next subsection, we describe the conditional sampling procedure.
Figure 3: A flow map of the Joint Diffusion Kalman Filter (JDKF) procedure.
@@ 3.4 Sampling $\psi_{t+1}$ Conditional on Information from $(x_{t+1},y_{t+1})$
4 Convergence of Averages and Robustness of Approximations
The works [shnitzer2020diffusion, shnitzer2022manifold] leverage the algorithm put forth in [singer2008nonlinear] to learn the eigenvalues of $\mathscr{L}$, to estimate the dynamics of the eigenfunctions, and to approximate the observations using a linear lift. The Kalman filtering framework of [shnitzer2020diffusion] is the motivation for the Joint Diffusion Kalman Filter and conditional sampling procedure developed in Subsections 3.3 and 3.4. However, there are some mathematical questions that need to be addressed to make this procedure rigorous. Under the assumptions in Subsection 3.1, we show in this section that it is admissible to learn the eigenvalues and the lift from time series data. In particular, a Poincaré Inequality (Assumption 3) implies that initialization errors are forgotten, and time averages converge (see Subsection 4.1). In Subsection 4.2, we bound the approximation error of the eigenfunctions of $\mathscr{L}$ by the diffusion map eigenvectors using a concentration inequality. Furthermore, in Subsection 4.3 we quantify the error of approximating the eigenfunction dynamics using linear diffusions, demonstrating robustness.
$S_{N}=\frac{1}{N+1}\sum_{i=0}^{N}V(\theta(t_{i})).$ (4.1)
For instance, in the setting of [singer_graph_2006], $(\theta(t_{i}))_{i=0}^{N}$ are independent uniform samples from a compact Riemannian manifold $\mathcal{M}$, and the graph Laplacian is shown to approximate the Laplace–Beltrami operator $\Delta_{\mathcal{M}}$, that is, the following holds for smooth $f$ and a point $\theta(t_{i})$ fixed:
In our setting, the summands in (4.1) arise not from independent samples, but from a sampled Langevin diffusion, which yields a Markov chain. In this case, the convergence of $S_{N}$ to the expectation $\mathbb{E}_{\mu}[V]$ as $N\to\infty$ is an ergodic property, which is guaranteed for $V\in\mathbb{L}^{2}(\mu)$ under our assumption of a Poincaré inequality or (equivalently) a spectral gap. If the Langevin diffusion is started in equilibrium, i.e. $\theta_{0}\sim\mu$, then $(\theta(t_{i}))_{i=0}^{N}$ is a stationary sequence and $\mathbb{E}_{\theta_{0}\sim\mu}[S_{N}]=\mathbb{E}_{\mu}[V]$ for any $N\geq 0$. In this case, the fluctuations $\sqrt{N}S_{N}$ converge as $N\to\infty$ to a Gaussian distribution by the celebrated Kipnis–Varadhan Central Limit Theorem [kipnis_central_1986]. A fast rate of convergence to equilibrium for Markov chains is a desirable property in the context of Markov chain Monte Carlo algorithms, and many central limit theorem results are collected in the survey paper [roberts_general_2004]. For continuous time averages, we refer the reader to [cattiaux_central_2012], where the trend to equilibrium is studied under weaker assumptions than ours (for instance, some laws with sub-exponential tails are shown to be ergodic). In general, the rate of convergence depends both on the integrability of $V$ and the weight of the tails of $\mu$. To fix notation, throughout the remainder of this section we consider a constant fixed step size $\Delta t=t_{i}-t_{i-1}$ and set $t_{0}=0$.
4.1 Short-Term Equilibration and Long-Term Convergence
where we have used Assumption 3 for the inequality. The differential form of Grönwall’s Inequality yields $\|P_{t}f\|_{\mathbb{L}^{2}(\mu)}^{2}\leq\exp(-2t/C)\|f\|_{\mathbb{L}^{2}(\mu)}^{2}$. Hence, $P_{t}f$ converges to equilibrium in $\mathbb{L}^{2}$ exponentially fast. In fact, this property is equivalent to a Poincaré Inequality (see for instance [bakry_analysis_2014, Theorem 4.2.5]). In the next lemma, we exploit this fact to bound the deviation of observables from their values at equilibrium when $\theta_{0}$ is initialized at some $\mu_{0}$ which is absolutely continuous with respect to the equilibrium measure $\mu$. In practice, this error can be diminished by discarding an initial burn-in period (as we have done in Subsection 5.2). Similar results hold in other $\mathbb{L}^{p}$-spaces by Riesz–Thorin Interpolation (see, for instance, [rudolf_explicit_2012, Proposition 3.17] and the application to MCMC methods in [fan_hoeffdings_2021, Theorem 12]), but we restrict to the $\mathbb{L}^{2}$ case to reduce notation.
Lemma 4.1.
Proof.
We have that
$\mathbb{E}_{\theta_{0}\sim\mu_{0}}\left[f(\theta_{t})\right]$ $=\int_{\mathbb{R}^{d}}f(w)\left(1+\frac{\mathrm{d}\mu_{t}}{\mathrm{d}\mu}(w)-1\right)\mu(\mathrm{d}w)$
$\leq\mathbb{E}_{\mu}[f]+\left\|\frac{\mathrm{d}\mu_{t}}{\mathrm{d}\mu}-1\right\|_{\mathbb{L}^{2}(\mu)}\left\|f\right\|_{\mathbb{L}^{2}(\mu)}$
$\leq\mathbb{E}_{\mu}[f]+\exp\left(-\lambda_{1}t\right)\left\|\frac{\mathrm{d}\mu_{0}}{\mathrm{d}\mu}-1\right\|_{\mathbb{L}^{2}(\mu)}\left\|f\right\|_{\mathbb{L}^{2}(\mu)}$
where we have used Hölder’s Inequality for the first inequality, and the $\mathbb{L}^{2}$-exponential convergence of $P_{t}$ for the second (recall that the spectral gap and first non-zero eigenvalue of $-\mathscr{L}$ satisfy $C=1/\lambda_{1}$). ∎
The first sum concerns the approximation of the eigenfunctions of $\mathscr{L}$ by diffusion coordinates, which is addressed in the next subsection. Focusing on the second term, we shall apply the Kipnis–Varadhan Central Limit Theorem [kipnis_central_1986]. We state the result with $\theta_{0}\sim\mu$; however, under suitable assumptions on $F^{j}\varphi_{k}$ similar results hold even when the Markov chain is started from a deterministic initial condition (see, for instance, [derriennic_central_2001, derriennic_central_2003, cuny_central_2012, cattiaux_central_2012]).
Proposition 4.2 ([kipnis_central_1986]).
holds, where $\sigma^{2}$ satisfies
and $\rho_{F^{j}\varphi_{k}}$ is the spectral measure of $F^{j}\varphi_{k}$.
4.2 Error Analysis of the Graph Laplacian
The ADM approach of [singer2008nonlinear] relies on the approximation of the Fokker–Planck operator by the graph Laplacian. Suppose $(\theta(t_{i}))_{i=0}^{N}$ are i.i.d. samples from $\mu$, let $f\in\mathbb{L}^{2}(\mu)$, and set $z(t_{i})=H(\theta_{i})$. Then an analogous argument to (4.2) (see [singer2008nonlinear, Equation (30)] along with extra accounting for a factor of 2 from differences in some definitions given there) yields
$\sum_{j=0}^{N}\mathbf{P}_{ij}f(z(t_{j}))\to\frac{\mathbb{E}_{\mu}\left[\exp\left(-\frac{\|\theta-\theta(t_{i})\|^{2}}{2\varepsilon}\right)f\circ H(\theta)\right]}{\mathbb{E}_{\mu}\left[\exp\left(-\frac{\|\theta-\theta_{j}\|^{2}}{2\varepsilon}\right)\right]}=f\circ H(\theta(t_{i}))+\varepsilon\mathscr{L}(f\circ H)(\theta(t_{i}))+O(\varepsilon^{2})$
Setting $w=\theta(t_{i})$, we have that
Theorem 4.3.
where $C$ depends on $\|u\|_{\infty}$.
Remark 4.4.
Proof.
We adapt the argument from [singer_graph_2006, Section 3] to samples $(\theta(t_{i}))_{i=0}^{N}$ from the Langevin diffusion (3.1). In this case, $(\theta(t_{i}))_{i=0}^{N}$ forms a reversible Markov chain with spectral gap $1-\exp(-\lambda_{1}\Delta t)$. We will show an upper bound for the right tail in the desired inequality. The left tail follows similarly and the statement of the theorem holds by a union bound. If we let
then we see that
Note that $\mathbb{E}_{\mu}[V]=0$ and $V$ is bounded in absolute value by some constant $C$ which depends on $\|u\|_{\infty}$. We seek to apply a concentration inequality which will capture the second moment of $V$ in the variance proxy. The version of Bernstein’s Inequality for Markov chains given in [paulin_concentration_2015, Theorem 3.3] gives
A lengthy computation (which is relegated to Appendix C) shows that
and the conclusion of the theorem follows. ∎
where $B^{i}$ is a Brownian motion, and for $i\neq j$ we have
Corollary 4.5.
Proof.
For $u\geq 0$, let $\mu_{u}$ denote the law of $\theta_{u}$. We have that
In [shnitzer2020diffusion], the gradients $\|\nabla\varphi_{i}(\theta_{t})\|_{2}$, for $i=1,\dots,\ell$ are assumed to be constant, so that the SDE for $\varphi_{i}(\theta_{t})$ is linear-Gaussian and a Kalman filtering framework may be applied. As another corollary of Lemma 4.1, we show next that the approximation error between the contractive dynamics (4.4) and a linear SDE can remain small. For $i=1,\dots,\ell$, let $\xi^{1},\dots,\xi^{\ell}$ satisfy
Corollary 4.6.
where
Remark 4.7.
Proof.
By Itô’s formula, the squared error at time $t\geq 0$ satisfies
For $u\geq 0$, let $\mu_{u}$ denote the law of $\theta_{u}$. By Lemma 4.1, we have
The final result follows from integrating in time, substituting into (4.7), and applying the integral form of Grönwall’s Inequality. ∎
5 Model Validation

### v2

@@ 5 Joint Diffusion Kalman Filter
Assuming that the gradients in (3.8) are constant, [36] define a state-space model, dubbed the Diffusion Kalman Filter (DKF), for the case where the data arises from a single data set $(x(t_{i}))_{i=0}^{N}$ (see Appendix A). We now modify DKF for a supervised learning framework to incorporate information in the series of the dependent variables $y(t_{i})$, which depend jointly with $x(t_{i})$ on $\theta(t_{i})$. If one were to use the estimated diffusion coordinates from DKF as developed in [36], a simple OLS regression of the dependent variables on these coordinates will show that most of the explanatory power of the original covariates $x(t_{i})$ for the dependent variables is lost (see, for instance, Appendix C for the regression of individual stock returns on a high-dimensional set of macroeconomic covariates). Thus, it is important to extend DKF to also account for the variation in the dependent variable in order for the diffusion coordinates to preserve explanatory power. We call this approach the Joint Diffusion Kalman Filter (JDKF).
Our framework relies on constructing dynamic diffusion coordinate embeddings of the observations that serve as proxies for the dynamics of $\theta$, combining state space modeling with diffusion maps. Commonly, dimensionality reduction methods, when applied to time series data, do not explicitly consider the temporal component in the data. This can be limiting in many applications, for instance, when dealing with high-dimensional financial market data where a portfolio manager (PM) might want to study how shocks to a given market factor propagate through time. By integrating Kalman filtering techniques with diffusion maps and extending the framework first introduced in [36], we construct dynamic diffusion coordinate embeddings which explicitly factor in the temporal component of the data. We construct our state space equations by leveraging a linear lifting operator that takes diffusion map embeddings back into the original measurement space.
Since we approximate the intrinsic dynamics with a linear SDE in terms of the diffusion coordinate embeddings (which we rigorously justify in Section 4.3), our framework does not require us to discover the explicit joint dynamics of the observations $(x(t),y(t))$. Leveraging linear state-space equations, we devise a procedure to sample dynamic embeddings while conditioning on an information set defined in the original measurement space. It is well known that dimensionality reduction techniques give rise to the challenge of interpretability of the embeddings. In a framework such as ours that takes place completely in the embedding space, it is hence desirable to define fully interpretable conditioning events in the original measurement space. To define such a sampling procedure, we leverage the power of the linear lifting operator in a three step process: (1) lift statistics from the embedding space to the original measurement space, (2) condition on the desired information set in the covariate space and obtain the conditional distribution in terms of $(x(t),y(t))$, (3) restrict back to the embedding space with a pseudoinverse of the lifting operator. Our sampling procedure simultaneously learns the intrinsic geometry and dynamics from the data, and then generates sample paths for the diffusion coordinate embeddings and by extension of $(x(t),y(t))$. Of course, it is known that dimensionality reduction techniques do not necessarily preserve statistics in the data, hence approximating statistics in the measurement space with their lifted counterparts incurs some error. Our robustness results presented in Section 4 provide theoretical support for the applicability of our method.
To formalize the Joint Diffusion Kalman Filter described above, we now specialize to the supervised learning framework and adopt the dynamical system setting from Subsection 3.1:
@@
Following the embedding procedure in Subsection 3.2, the diffusion coordinates $\psi(t_{i})$ are computed jointly on samples of $(X,Y)$, denoted $(x(t_{i}),y(t_{i}))_{i=0}^{N}$, and $x$ and $y$ can be similarly reconstructed by the linear lifts:
@@
where $\mathbf{H}_{i,j}^{x}=\langle x_{i},\psi_{j}\rangle_{\text{data}}$ and $\mathbf{H}_{i,j}^{y}=\langle y_{i},\psi_{j}\rangle_{\text{data}}$. Let $\Lambda$ denote the transition matrix for the linear dynamics containing the mapped eigenvalues $\varepsilon^{-1}\log\kappa_{1},\dots-\varepsilon^{-1}\log\kappa_{\ell}$} on its diagonal.
@@
5.1 E-Step: Kalman Filter with Joint Observations
The E-step estimates the latent diffusion coordinates $\psi_{t}$ and their covariance $P_{t|t}$ using a Kalman filter that processes joint observations from both covariates and responses. We define the joint observation vector:
$z_{t}=\begin{bmatrix}x_{t}\\ y_{t}\end{bmatrix}=\begin{bmatrix}\mathbf{H}^{x}\\ \mathbf{H}^{y}\end{bmatrix}\psi_{t}+\begin{bmatrix}v_{x,t}\\ v_{y,t}\end{bmatrix}=\mathbf{H}^{z}\psi_{t}+v_{t}$
where $\mathbf{H}^{z}=[\mathbf{H}^{x};\mathbf{H}^{y}]$ and the joint noise covariance is:
$\mathbf{R}=\begin{bmatrix}\mathbf{R}_{x}&\mathbf{R}_{xy}\\ \mathbf{R}^{\top}_{xy}&\mathbf{R}_{y}\end{bmatrix}$.
The crucial observation is that by filtering on the stacked observation vector $z_{t}$, the Kalman gain optimally weights information from both covariates and responses when updating the latent state estimate. This contrasts with DKF [36], which processes only a single observation stream and thus may lose explanatory power for the responses.
Running the forward Kalman filter yields filtered estimates $\hat{\psi}_{t|t}$ and covariances $P_{t|t}$ for $t=1,\ldots,T$. We then apply the Rauch–Tung–Striebel (RTS) smoother [33] in a backward pass to obtain smoothed estimates $\hat{\psi}_{t|T}$ and $P_{t|T}$ that incorporate information from the entire observation sequence. The standard Kalman filtering and RTS smoothing equations are provided in Appendix B for completeness.
In the linear multi-factor case where $\mathbf{H}^{y}=\mathbf{B}\mathbf{H}^{x}$, this formulation naturally accommodates the correlation structure induced by the factor loadings $\mathbf{B}$.
5.2 M-Step: Parameter Updates
The M-step updates the observation noise covariances based on the smoothed state estimates from the E-step. Given $\hat{\psi}_{t|T}$ for $t=1,\ldots,T$, we update:
  • $\mathbf{R}_{x}$: covariance of residuals $x_{t}-\mathbf{H}^{x}\hat{\psi}_{t|T}$
  • $\mathbf{R}_{y}$: covariance of residuals $y_{t}-\mathbf{H}^{y}\hat{\psi}_{t|T}$
  • $\mathbf{R}_{xy}$: cross-covariance of these residuals
@@ 5.2 M-Step: Parameter Updates
In the linear multi-factor case, we additionally estimate the factor loading matrix $\mathbf{B}$ via least squares regression on the lifted coordinates:
$\mathbf{B}=\left(\sum_{t=1}^{T}y_{t}(\mathbf{H}^{x}\hat{\psi}_{t|T})^{\top}\right)\left(\sum_{t=1}^{T}\mathbf{H}^{x}\hat{\psi}_{t|T}(\mathbf{H}^{x}\hat{\psi}_{t|T})^{\top}\right)^{-1}.$
This update is consistent with the observation equation $y_{t}=\mathbf{B}\mathbf{H}^{x}\psi_{t}+v_{y,t}$ and ensures that the estimated loadings reflect the relationship between responses and the lifted covariate embeddings.
The complete parameter update formulas are standard and provided in Appendix B. The key distinction from classical EM for state-space models is that our observation matrices $\mathbf{H}^{x},\mathbf{H}^{y}$ are learned from data via diffusion maps rather than specified a priori, and the transition matrix $\mathbf{A}=I-\Lambda$ is determined by the diffusion map eigenvalues rather than estimated from the data.
5.3 EM Algorithm
@@
We have added an extra layer with the measurement equation for the dependent variables to maintain the explanatory power in the embeddings and taken this into account via the modified state-space. JDKF thus gives us diffusion coordinates that propagate linearly and preserve much of the original explanatory power from the full high-dimensional set of covariates. Algorithm 1 summarizes the EM procedure. Figure 3 displays a broad flow map of the JDKF procedure. In the next subsection, we describe the conditional sampling procedure.
Figure 3: A flow map of the Joint Diffusion Kalman Filter (JDKF) procedure. Starting from observations $z(t_{i})$, we apply anisotropic diffusion maps to obtain embeddings $\psi(t_{i})$ that capture the manifold geometry. We propagate a linear approximation of the diffusion coordinates forward in time. We predict coordinates using a Kalman filter and lift back to the measurement space using the linear operator $\mathbf{H}$.
@@ 5.4 Sampling $\psi_{t+1}$ Conditional on Information from $(x_{t+1},y_{t+1})$
6 Application to Financial Stress Testing